OpenStudy (anonymous):

Simplify the expression. 2(-4 - 3)/4 + 3 x 2 A. 3.25 B. 2.5 C. 0.5

jhonyy9 (jhonyy9):

-7/2 +3x2 =0 6x2=7 x2=7/6 x=sqrt(7/6)

OpenStudy (precal):

2(-4-3)/4+3x2 2(-7)/4+3x2 -14/4+3x2 -3.5+6 2.5 I am assuming that it was 3 times 2 at the end. Use order of operation PEMDAS